Following is a trial balance (in thousands) showing the accounts of Beta Hospital at December 31, 2019. Using the information from the trial balance, answer the following questions. Show the calculations used.

  1. What is number of days’ cash on hand?
  2. What is the number of days of revenue in net patient accounts receivable?
  3. How much is the excess of revenues over expenses?
  4. What is the operating margin?
  5. What is the long-term debt-to-equity ratio at year-end?
  6. What is the interest coverage (times interest earned)?
Debits Credits
Cash $5,000
Patient accounts receivable 8,000
Allowance for bad debts $1,000
Food and medicines inventory 2,000
Buildings 40,000
Accumulated depreciation, buildings 8,000
Equipment 20,000
Accumulated depreciation, equipment 10,000
Land 5,000
Accounts payable 6,000
Accrued interest payable 200
Current portion of bonds payable 1,500
Long-term bonds payable 34,000
Net position, January 1, 2019 17,300
Patient service revenues (net) 36,000
Patient care expense 20,000
Dietary expense 2,000
General and administrative expense 6,000
Bad debts expense 1,000
Depreciation, buildings and equipment 4,000
Interest expense 1,000
Totals $114,000 $114,000
Answer #1
Requirement: Formula Result
Days cash on hand cash*365/Revenues 5000*365/36000 = 50.70 days
Days in Net Patient AR Net AR*365/Revenues 7000*365/36000 = 70.97 days
Excess of revenues over expenses Total revenues - Total expenses 36000 - (20000+2000+6000+1000+4000+1000)=$2000
Operating Margin Operating Income/Total Revenues 2000/36000 = 5.56%
Long term debt to Equity ratio LTD / EQUITY 34000/(17300+2000) = 1.76
Interest coverage Interest paid / (Total revenues - Direct expenses) 1000/(36000-(20000+2000)) = 0.07